In a science laboratory, a group of scientists would like to test the reaction of argon and ethane (C2H6) vapor. A 2.00 L cylindrical container is filled with 3.50 g of argon at 60 ◦C. Ethane vapor is the added to the same container until the total pressure reached to 2.50 atm. Calculate the mass of ethane that was added in the container?


1
Expert's answer
2021-12-21T14:46:01-0500

Molar mass of Argon= 39.948

Molar mass of Ethane= 30.0694

R= 0.08206 L atm/ mol K

60°C = 333K


Partial pressure of Argon:


PV =n RT

"n_{ar}=\frac{3.5}{39.948}= 0.08761\\P_{ar}= \frac{0.08761×0.08206×333}{2}\\=1.1970"

Partial pressure of Ethane


"2.5-1.1970\\=1.3030"


"\frac{n_{ar}}{n_{ethane}}=\frac{P_{Ar}}{P_{ethane}}"


"\frac{0.08761}{n_{ethane}}=\frac{1.1970}{1.3030}"


"n_{ethane}=0.09537"


Mass of ethane "=0.09537×30.0694\\"


"=2.868g"